# A discrete approach to Dirichlet L-functions via spectral models: special values and zeros

### v1

@@ Abstract
We study Dirichlet $L$-functions via discrete analogs $L_{n}$ arising from the spectral theory of cyclic graphs as $n\rightarrow\infty$. Using a refined Euler-Maclaurin asymptotic expansion due to Sidi, together with an independent polynomiality property of these finite spectral sums at integers, we obtain exact special-value formulas, even starting at $n=1$. This yields new expressions for certain trigonometric sums of interest in physics, and recovers, by a different method, the striking formulas of Xie, Zhao, and Zhao. It also gives infinite families of recursion relations among special values of the Riemann zeta function and of Dirichlet $L$-functions.
Concerning zeros, we prove that, for odd primitive characters, a natural asymptotic functional equation for the discrete functions $L_{n}$ is equivalent to the Generalized Riemann Hypothesis for the corresponding Dirichlet $L$-function. We also provide some remarks about the non-existence of possible real zeros.
@@ 1 Introduction
Dirichlet $L$-functions, defined for characters $\chi$ by the meromorphic continuation of
@@
play a central role in analytic number theory. Our perspective in this paper is to view these functions as spectral objects arising from the usual Laplacian, and to approximate them by quantities defined in a completely parallel way from the combinatorial Laplacian. This continues the approach set forth in [CJK10, FK17, F16, MV23, XZZ24], and the focus of our paper is special values and zeros of $L(s,\chi)$, along with exact evaluations of certain trigonometric sums that appear in other areas.
It should be said that the closest graph analog of Selberg’s zeta function is instead the Ihara zeta functions, for example since they both are Laplace transforms of heat kernels, see [CJK15]. The Ihara zeta function for general graphs was suggested by Serre in the 1970s [S02, p. IX] after reading Ihara’s work. It was developed by Sunada and others into an important chapter of spectral graph theory [Te10]. These zetas (and the $L$-functions of Stark-Terras [ST96]) have analogs with number theoretic zeta functions. In contrast, the functions associated to graph spectra discussed here have more direct connections. Selberg’s zeta function, with its Euler product and trace formula, has features analogous to classical number theory. But in addition, for the modular surface, the Riemann zeta function actually appears in the functional equation.
$\zeta_{\mathbb{Z}}(s)=\frac{1}{\Gamma(s)}\int_{0}^{\infty}e^{-2t}I_{0}(2t)t^{s}\frac{dt}{t}=\frac{1}{2^{2s}\pi^{1/2}}\frac{\Gamma(1/2-s)}{\Gamma(1-s)}=\binom{-2s}{-s}$.
It has its share of analogs, for example a functional equation of the usual form $\xi_{\mathbb{Z}}(1-s)=\xi_{\mathbb{Z}}(s)$ [FK17]. Moreover, special values of $\zeta_{\mathbb{Z}}$ can be used to give the formula of the volumes of the round spheres in a form similar to the Minkowski-Siegel volume formulas for spaces like $\mathrm{SL}_{n}(\mathbb{R})/\mathrm{SL}_{n}(\mathbb{Z})$ [KP23]. But $\zeta_{\mathbb{Z}}(s)$ also actually appears in the functional equation of Eisenstein series [Se56, (3.7),(3.9)], and of Selberg’s zeta function [Se56, p. 79] and [Se54, §10]. Thus, in the case of the modular surface, both $\zeta(s)$ and $\zeta_{\mathbb{Z}}(s)$ appear in the fudge factors for the Selberg zeta function (cf. [JL11, p. 300]). This may not be surprising since both $\mathbb{R}/\mathbb{Z}$ and $\mathbb{Z}$ can be thought related to the cusp. The graph zeta function $\zeta_{\mathbb{Z}}(s)$ is also present in the Gindikin-Karpelevich formulas [K20], and as Sarnak pointed out to us, it is suggestive that the special values at s=ns=-n are exactly the integers exploited by Chebyshev in his work on primes.
An infinite sum, as the Dirichlet series displayed above, is by definition the limit of the partial sums. Here we take instead limits of finite sums with a much richer structure: they are defined entirely analogously as Mellin transforms of heat kernels, like in Riemann’s memoir or as emphasized by Jorgenson-Lang [JL11]. Thanks to this structure, these sums have independent interests in various physics and mathematics contexts, see [K20, JKS24] for references. The special value $\zeta^{\prime}(0)$ is the negative of the logarithm of the determinant of the Laplacian, which counts the number of (rooted) spanning trees and therefore has a long history. It is a graph invariant with exceptionally wide relevance, for example to the study of water systems [BR18].
@@
Friedli [F16] studied the following discrete analog of Dirichlet $L$-functions:
@@
Asymptotics.
Similarly to [FK17] in the case of the Riemann zeta function, see Theorem 3.1 below, and [F16] for even characters, see Theorem 3.2 below, we use a modified Euler-Maclaurin formula that allows for endpoint singularities, developed by Sidi in [Sid04] and [Sid12] in the context of numerical analysis, to establish:
@@
Let $\chi$ be an odd Dirichlet character of modulus $q>1$. Then, for any fixed $s\in\mathbb{C}$ and any integer $m\geq 0$, the following asymptotic expansion holds as $n\to\infty$:
$L_{n}(s,\chi)=2\sum_{k=0}^{m}b_{k}(s)L(1+2(s-k),\chi)\left(\frac{qn}{2\pi}\right)^{1+2(s-k)}+\mathcal{O}(n^{1+2(s-(m+1))}),$
@@
Special values.
The fact that limits of certain trigonometric sums provide a method to evaluate Euler’s special values for $\zeta(s)$ has been observed in several places, for example by Apostol [Ap73] or Witten [Wi91, p. 177-178]. A remarkable discovery of Xie-Zhao-Zhao [XZZ24] is that there is in fact no need to take a limit. The finite cyclic graph approximates the circle, like in Archimedes, but in contrast for special values there is no need for any limits. In a precise sense, see below, already the graph on one vertex and one loop, or the graph on two vertices with a double edge, know $\pi^{2}/6$.
Moreover, the authors of [XZZ24] extends this to special values of Dirichlet $L$-functions, to obtain closed form expressions in terms of the discrete $L$-functions considered here. The special case of Riemann zeta functions actually already appears in a paper by Zagier [Za96, Theorem 1.iii], although via different methods. It seems unlikely that Zagier thought of it as a graph zeta function, nor did he emphasize Riemann’s zeta function. Instead, he was concerned with Verlinde formulas, see [K20] for this connection to zeta functions.
In the present paper, our method appears flexible and conceptually attractive. It is the combination of a refined Euler–Maclaurin–Sidi asymptotics with the structural polynomiality of the spectral zeta functions. This gives that an asymptotic expansion that terminates because of trivial zeros must in fact coincide with the exact finite $n$ identity. Curiously, this allows one to derive nontrivial closed formulas already at $n=1$ from an argument initially designed for $n\rightarrow\infty$.
This asymptotic-to-exact mechanism seems applicable in other settings where spectral data on finite graphs approximates a continuous model, and may be of independent interest.
We prove the following theorem, and refer to [JKS24] for previous results and pointers to the extensive literature about such sums. In particular, they appear in several physics contexts. First, let $s\mapsto a_{k}(s)$ be the polynomial functions of degree $k$, given by the Taylor expansion
@@
For any $0<\theta<1$ and integers $n,p\geq 1$, the following holds:
$\sum_{j=0}^{n-1}\sin^{-2p}\left(\frac{\pi(j+\theta)}{n}\right)=2^{2p}\sum_{k=0}^{p}a_{p-k}(p)\zeta_{\mathbb{R}/\mathbb{Z}}(k,\theta)n^{2k},$ (1)
@@
$\sum_{j=0}^{n-1}\cot\left(\frac{\pi(j+\theta)}{n}\right)\left[\sin\left(\frac{\pi(j+\theta)}{n}\right)\right]^{-2p}=-\frac{2^{2p}}{2\pi p}\sum_{k=0}^{p}a_{p-k}(p)\,\partial_{\theta}\zeta_{\mathbb{R}/\mathbb{Z}}(k,\theta)n^{2k+1}.$ (2)
The $\zeta_{\mathbb{R}/\mathbb{Z}}$ is the spectral zeta function associated to a constant bundle over the circle depending on $\theta$. As such, it is a version of the Hurwitz zeta function, and the corresponding special values at integers involve Bernoulli polynomials and polygamma functions, see section 2.1 below. In particular, applying the theorem with $n=1$ recovers part of a well-known reflection relation for the polygamma functions.
The formulas in Theorem 1.2, and in Theorem 1.4 below, can be viewed as giving infinite families of recursion relations between special values. Here is the case of the Riemann zeta function:
@@
For any integers $n,p\geq 1$, the following holds:
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(p)=\sum_{k=0}^{p}a_{p-k}(p)\zeta_{\mathbb{R}/\mathbb{Z}}(k)n^{2k}.$
In particular for $n=1$,
$\zeta(2p)=-\sum_{k=0}^{p-1}a_{p-k}(p)(2\pi)^{2p-2k}\zeta(2k),$
and for $n=2$,
$\zeta(2p)=2^{-2p-1}(2\pi)^{2p}-\sum_{k=0}^{p-1}a_{p-k}(p)2^{-2k}(2\pi)^{2p-2k}\zeta(2k).$
These formulas generate infinite families of linear recursion relations among the values $\zeta(2p)$, extending earlier more isolated examples in the literature, see [Me17] for a recent study. In view of Euler’s evaluation, these formulas also translate into linear recursion relation between Bernoulli numbers. Shifting the viewpoint, they can also be viewed as Faulhaber-type formulas for trigonometric sums.
@@
Let $\chi$ be an even Dirichlet character of modulus $q$ (not necessarily primitive). Then for any integers $n,p\geq 1$,
$L_{n}(p,\chi)=2\sum_{k=0}^{p}a_{p-k}(p)L(2k,\chi)\left(\frac{qn}{2\pi}\right)^{2k}.$
Let $\chi$ be an odd Dirichlet character of modulus $q$ (not necessarily primitive). Then for any integers $n,p\geq 1$,
$L_{n}(p,\chi)=2\sum_{k=0}^{p}\frac{2k}{p}a_{p-k}(p)\,L(2k+1,\chi)\left(\frac{qn}{2\pi}\right)^{2k+1}.$
In the case $n=1$ these formulas were found by Xie, Zhao, and Zhao in [XZZ24] as mentioned above. They also displayed the corresponding formulas expressing the special values of the Dirichlet $L$-functions in terms of the discrete ones.
Zeros of Dirichlet $L$-functions.
The asymptotics in Theorem 1.1 lead to a statement about possible Siegel zeros:
Corollary 1.5 (of Theorem 1.1).
@@
Not only is there a connection via the sign, or the vanishing, of $L_{n}(s,\chi)$, to zeros of $L(s,\chi)$, but there is a more surprising connection. Indeed, the Riemann Hypothesis (RH) in question is equivalent to a graph zeta- or $L$-function asymptotic functional equations. It has been established in the following cases:
@@
In each case, the form and algebra of the proofs differ. Therefore it does not seem to be a one-time accident. In all cases, it relies on nontrivial monotonicity results of Matiyasevich-Saidak-Zvengrowski [MSZ14]. To this list we can now add the odd character version, which completes the picture in the $1$-dimensional case. Let
@@
Remark: In view of the standard factorization of the above-mentioned Dedekind zeta function, there are thus two graph reformulations of its RH: via cyclic graphs and via two-dimensional discrete tori.
@@ 2.3 Asymptotics
For any $s\neq\tfrac{1}{2},\tfrac{3}{2},\tfrac{5}{2},\ldots$, any $0<\theta<1$, and any $m\geq 0$, the following asymptotics hold as $n\to\infty$:
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(s,\theta)=n\zeta_{\mathbb{Z}}(s)+n^{2s}\left(\sum_{k=0}^{m}a_{k}(s)\zeta_{\mathbb{R}/\mathbb{Z}}(s-k,\theta)n^{-2k}+\mathcal{O}(n^{-2-2m})\right),$ (7)
@@
$\partial_{\theta}\zeta_{\mathbb{Z}/n\mathbb{Z}}(s,\theta)=n^{2s}\left(\sum_{k=0}^{m}a_{k}(s)\partial_{\theta}\zeta_{\mathbb{R}/\mathbb{Z}}(s-k,\theta)n^{-2k}+\mathcal{O}(n^{-2-2m})\right),$ (8)
@@
$\sum_{j=0}^{n-1}f_{s}\left(\frac{j+\theta}{n}\right)$ $=n\int_{0}^{1}f_{s}(x)\mathrm{d}x+\sum_{k=0}^{m}a_{k}(s)(\zeta(2(s-k),\theta)+\zeta(2(s-k),1-\theta))\left(\frac{2\pi}{n}\right)^{2(k-s)}$
$+\mathcal{O}(n^{2(s-m-1)}).$
@@
$+\underset{k\neq s}{\sum_{0\leq k\leq m}}b_{k}(s)(\zeta(1+2(s-k),\theta)-\zeta(1+2(s-k),1-\theta))\left(\frac{2\pi}{n}\right)^{2(k-s)-1}$
$+\mathcal{O}(n^{1+2(s-m-1)}),$
@@
$\sum_{j=0}^{n-1}g_{s}\left(\frac{j+\theta}{n}\right)$ $=-\frac{n}{2\pi s}\Bigg[\sum_{k=0}^{m}b_{k}(s)\frac{s}{2(s-k)}\partial_{\theta}\zeta_{\mathbb{R}/\mathbb{Z}}(s-k,\theta)n^{2(s-k)}$
$+\mathcal{O}(n^{2(s-m-1)})\Bigg],$

### v2

# A discrete approach to Dirichlet L-functions, their special values and zeros

@@ Abstract
We obtain new infinite families of identities among special values of Dirichlet $L$-functions using finite spectral sums. More precisely, we study Dirichlet $L$-functions via discrete analogues $L_{n}$ arising from the spectral theory of cyclic graphs as $n\rightarrow\infty$. Applying a refined Euler-Maclaurin asymptotic expansion due to Sidi, together with an independent polynomiality property of these finite spectral sums at integers, we obtain exact special-value formulas, even starting at $n=1$. This yields new expressions for certain trigonometric sums of interest in physics, and recovers, by a different method, the striking formulas of Xie, Zhao, and Zhao.
Concerning zeros, using the same asymptotic expansion, we prove that for odd primitive characters, an asymptotic functional equation relating $L_{n}(1-s,\overline{\chi})$ to $L_{n}(s,\chi)$ is equivalent to the Generalized Riemann Hypothesis for the corresponding Dirichlet $L$-function $L(s,\chi)$. We also provide some remarks about the non-existence of possible real zeros.
@@ 1 Introduction
Dirichlet $L$-functions, defined by the meromorphic continuation of
@@
play a central role in analytic number theory, through their Euler product, special values, functional equation, and zeros. In this paper we use a spectral viewpoint and asymptotics that simultaneously lead to infinite families of new recursive identities among special values (Corollaries 1.3 and 1.4), exact formulas for certain trigonometric sums (Theorem 1.2), and a reformulation of the Generalized Riemann Hypothesis in terms of a certain asymptotic functional equation (Theorem 1.6).
For example, generalizing the formulas of Xie-Zhao-Zhao [XZZ24, Theorem A], we show that for any even Dirichlet character $\chi$ mod $q$ and any integers $m,n>0$,
[…]
where $a_{k}(s)$ are universal, explicit polynomials of degree $k$ with rational coefficients. In the case of the Riemann zeta function $\zeta(s)$, the simplest identity of this type is
[…]
See Corollary 1.3 for an infinite family of such relations. A version of this corollary already appeared in a paper by Zagier on Verlinde formulas [Za96, Theorem 1.iii], and other types of recursive relations can be found in the works of Euler, Ramanujan and others [SD86, Me17].
Our method is to view $\zeta(s)$ and $L(s,\chi)$ as spectral objects arising from the usual Laplacian, and to approximate them by quantities defined in a completely parallel way from the combinatorial Laplacian. This continues the approach set forth in [CJK10, FK17, F16, MV23, XZZ24].
An infinite sum, as the Dirichlet series displayed above, is by definition the limit of the partial sums. Here we take instead limits of finite sums with a much richer structure: they are defined entirely analogously as Mellin transforms of heat kernels, like in Riemann’s memoir or as emphasized by Jorgenson-Lang [JL11]. Thanks to this structure, these sums have independent interests in various physics and mathematics contexts, see [K20, JKS24] for references. For example, the special value $\zeta_{X}^{\prime}(0)$ is the logarithm of the number of rooted spanning trees, which is an invariant with relevance from electrical circuit theory to the study of water systems [BR18].
@@
Friedli [F16] studied the following discrete analogue of Dirichlet $L$-functions:
@@
1.1 Asymptotics.
Similarly to [FK17] in the case of the Riemann zeta function, see Theorem 3.1 below, and by Friedli [F16] for even characters, see Theorem 3.2 below, we use a modified Euler-Maclaurin formula that allows for endpoint singularities, developed by Sidi in [Sid04] and [Sid12] in the context of numerical analysis. We get:
@@
Let $\chi$ be an odd Dirichlet character of modulus $q>1$. Then, for any fixed $s\in\mathbb{C}$ and any integer $l\geq 0$, the following asymptotic expansion holds as $n\to\infty$:
$L_{n}(s,\chi)=2\sum_{k=0}^{l}b_{k}(s)L(1+2(s-k),\chi)\left(\frac{qn}{2\pi}\right)^{1+2(s-k)}+\mathcal{O}(n^{1+2(s-(l+1))}),$
@@
1.2 Special values.
The fact that limits of certain trigonometric sums provide a method to evaluate Euler’s special values for $\zeta(s)$ has been observed in several places, for example by Apostol [Ap73] or Witten [Wi91, p. 177-178]. A remarkable discovery of Xie-Zhao-Zhao [XZZ24] is that there is in fact no need to take a limit. The finite cyclic graph approximates the circle, like in Archimedes, but in contrast, for special values there is no need for any limits. In a precise sense, already the graph on one vertex and one loop knows $\pi^{2}/6$ (from equation (1) above with $m=1$ and $\zeta(0)=-1/2$).
Our method is to combine a refined Euler–Maclaurin–Sidi asymptotics with the structural polynomiality of the spectral zeta functions. This gives that, an asymptotic expansion that terminates because of trivial zeros, must in fact coincide with the exact finite $n$ identity. Curiously, this allows one to derive nontrivial closed formulas already at $n=1$ from an argument initially designed for $n\rightarrow\infty$.
This asymptotic-to-exact mechanism seems applicable in other settings, and we think that the method will yield further applications.
We prove the following theorem, and refer to [JKS24] for previous results and pointers to the extensive literature about such sums. First, let $s\mapsto a_{k}(s)$ be the polynomial functions of degree $k$, given by the Taylor expansion
@@
For any $0<\theta<1$ and integers $n,m\geq 1$, the following holds:
$\sum_{j=0}^{n-1}\sin^{-2m}\left(\frac{\pi(j+\theta)}{n}\right)=2^{2m}\sum_{k=0}^{m}a_{m-k}(m)\zeta_{\mathbb{R}/\mathbb{Z}}(k,\theta)n^{2k},$ (2)
@@
$\sum_{j=0}^{n-1}\cot\left(\frac{\pi(j+\theta)}{n}\right)\left[\sin\left(\frac{\pi(j+\theta)}{n}\right)\right]^{-2m}=-\frac{2^{2m}}{2\pi m}\sum_{k=0}^{m}a_{m-k}(m)\,\partial_{\theta}\zeta_{\mathbb{R}/\mathbb{Z}}(k,\theta)n^{2k+1}.$ (3)
The $\zeta_{\mathbb{R}/\mathbb{Z}}(s,\theta)$ is the spectral zeta function associated to a certain bundle over the circle depending on $\theta$. As such, it is a version of the Hurwitz zeta function, and the corresponding special values at integers involve Bernoulli polynomials and polygamma functions, see section 2.1 below. In particular, applying the theorem with $n=1$ recovers part of a well-known reflection relation for the polygamma functions. The previous theorem can be extended to any integers $m\in\mathbb{Z}$. This is described in section 6. In particular, for $m\leq 0$,
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(-m,\theta)=n\zeta_{\mathbb{Z}}(-m),$ (4)
for all $0<\theta<1$, and all $n>m$. See the end of this introduction for the definition of $\zeta_{\mathbb{Z}}(-m)$ (with no dependence on $\theta$). Here is the case of the Riemann zeta function:
@@
For any integers $n,m\geq 1$, the following holds:
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(m)=\sum_{k=0}^{m}a_{m-k}(m)\zeta_{\mathbb{R}/\mathbb{Z}}(k)n^{2k}.$
Explicitly, the case $n=1$ is the formula (1) above, and in general for $n\geq 1$,
$\frac{(-1)^{m+1}}{4}\sum_{a=0}^{2m}n^{-a}\left(\begin{array}[]{c}2m+1\\ a+1\end{array}\right)\sum_{j=0}^{a+1}(-1)^{j}\left(\begin{array}[]{c}a+1\\ j\end{array}\right)\frac{a+1-2j}{a+1}\left(\begin{array}[]{c}m+jn+(a-1)/2\\ 2m+a\end{array}\right)$
[…]
The left hand side in the last formula comes from [CM99, BY02]. The formulas in Corollary 1.3 should be compared with those in Zagier’s [Za96]. It seems unlikely that he thought of the trigonometric sums as special values of graph zeta functions. See also [XZZ24, Theorem B]. There are also more isolated examples of such recursive relations in the literature, see [SD86, Me17] and references therein.
@@
Let $\chi$ be an even Dirichlet character of modulus $q$ (not necessarily primitive). Then for any integers $n\geq 1$ and $m\geq 0$,
$L_{n}(m,\chi)=2\sum_{k=0}^{m}a_{m-k}(m)L(2k,\chi)\left(\frac{qn}{2\pi}\right)^{2k}.$
Let $\chi$ be an odd Dirichlet character of modulus $q$ (not necessarily primitive). Then for any integers $n\geq 1$ and $m\geq 0$,
$L_{n}(m,\chi)=2\sum_{k=0}^{m}b_{m-k}(m)\,L(2k+1,\chi)\left(\frac{qn}{2\pi}\right)^{2k+1}.$
In the case $n=1$ these formulas were found by Xie, Zhao, and Zhao in [XZZ24]. They also displayed the corresponding formulas expressing the special values of the Dirichlet $L$-functions in terms of the discrete ones.
Special values at negative integers m<0-m<0 of discrete $L$ functions satisfy
$L_{n}(-m,\chi)=0,$
whenever $n>m$ (see the proof of Corollary 6.5). These values should not be dismissed as trivial. For example, when $\chi$ is real, even and non-principal the condition $L_{n}(-m,\chi)\geq 0$ for all even $n\geq 2$ and for any integers $m>0$ implies the absence of real zeros in the critical strip of the corresponding Dirichlet $L$-function.
The case $m=0$ for odd characters in Corollary 1.4 has a special interest from the Dirichlet class number formula and reads
$L(1,\chi)=\frac{\pi}{2}\frac{1}{qn}\sum_{j=1}^{qn-1}\chi(j)\cot\left(\frac{\pi j}{qn}\right),$
for all $n\geq 1$. By trigonometric manipulations, this formula reduces to the classical cotangent formula (the case $n=1$) for $L(1,\chi)$.
1.3 Zeros of Dirichlet $L$-functions.
The asymptotics in Theorem 1.1 that was used to derive identities among special values, provide, at the same time, information about zeros. We begin by:
Corollary 1.5.
@@
Not only is there a connection via the sign, or the vanishing, of $L_{n}(s,\chi)$, to zeros of $L(s,\chi)$, but there is a more surprising connection. Indeed, the Riemann Hypothesis (RH) in question is equivalent to graph zeta- or $L$-function asymptotic functional equations. It has been established in the following cases:
@@
In each case, the form and algebra of the proofs differ. In all cases, it relies on nontrivial monotonicity results of Matiyasevich-Saidak-Zvengrowski [MSZ14]. To this list we can now add the odd character version, which completes the picture in the $1$-dimensional case. Let
@@
In view of the standard factorization of the above-mentioned Dedekind zeta function, there are thus two graph reformulations of its RH: via cyclic graphs and via two-dimensional discrete tori.
@@
1.4 Context.
It should be said that the closest graph analogue of Selberg’s zeta function is instead the Ihara zeta functions, for example since they both are Laplace transforms of heat kernels, see [CJK15]. The Ihara zeta function for general graphs was suggested by Serre in the 1970s [S02, p. IX] after reading Ihara’s work. It was developed by Sunada and others into an important chapter of spectral graph theory [Te10]. These zetas (and the $L$-functions of Stark-Terras [ST96]) have analogues with number theoretic zeta functions. In contrast, the functions associated to graph spectra discussed here have more direct connections. Selberg’s zeta function, with its Euler product and trace formula, has features analogous to classical number theory. But in addition, for the modular surface, the Riemann zeta function actually appears in the functional equation.
$\zeta_{\mathbb{Z}}(s)=\frac{1}{\Gamma(s)}\int_{0}^{\infty}e^{-2t}I_{0}(2t)t^{s}\frac{dt}{t}=\frac{1}{2^{2s}\pi^{1/2}}\frac{\Gamma(1/2-s)}{\Gamma(1-s)}=\binom{-2s}{-s}=\prod_{k=1}^{\infty}\frac{(1-s/k)}{(1-s/(k+1/2))},$
see [FK17, KP23]. It has its share of analogues, for example a functional equation of the usual form $\xi_{\mathbb{Z}}(1-s)=\xi_{\mathbb{Z}}(s)$. Moreover, special values of $\zeta_{\mathbb{Z}}$ can be used to give the formula of the volumes of the round spheres in a form similar to the Minkowski-Siegel volume formulas for spaces like $\mathrm{SL}_{n}(\mathbb{R})/\mathrm{SL}_{n}(\mathbb{Z})$ [KP23]. But $\zeta_{\mathbb{Z}}(s)$ also actually appears in the functional equation of Eisenstein series [Se56, (3.7),(3.9)], and of Selberg’s zeta function [Se56, p. 79] and [Se54, §10]. Thus, in the case of the modular surface, both $\zeta(s)$ and $\zeta_{\mathbb{Z}}(s)$ appear in the fudge factors for the Selberg zeta function (cf. [JL11, p. 300]). This may not be surprising since both $\mathbb{R}/\mathbb{Z}$ and $\mathbb{Z}$ can be thought related to the cusp. The graph zeta function $\zeta_{\mathbb{Z}}(s)$ is also present in the Gindikin-Karpelevich formulas [K20] and in the significant scaling-limit context of [ABS25].
Moreover, as Sarnak pointed out to us, it is suggestive that the special values at s=ms=-m are exactly the integers exploited by Chebyshev in his work on primes. Thus the graph special values, $\zeta_{\mathbb{Z}}(-m)$ and $L_{n}(-m,\chi)$, all have implications for the distribution of primes.
@@ 2.3 Asymptotics
For any $s\neq\tfrac{1}{2},\tfrac{3}{2},\tfrac{5}{2},\ldots$, any $0<\theta<1$, and any $l\geq 0$, the following asymptotics hold as $n\to\infty$:
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(s,\theta)=n\zeta_{\mathbb{Z}}(s)+n^{2s}\left(\sum_{k=0}^{l}a_{k}(s)\zeta_{\mathbb{R}/\mathbb{Z}}(s-k,\theta)n^{-2k}+\mathcal{O}(n^{-2-2l})\right),$ (9)
@@
$\partial_{\theta}\zeta_{\mathbb{Z}/n\mathbb{Z}}(s,\theta)=n^{2s}\left(\sum_{k=0}^{l}a_{k}(s)\partial_{\theta}\zeta_{\mathbb{R}/\mathbb{Z}}(s-k,\theta)n^{-2k}+\mathcal{O}(n^{-2-2l})\right),$ (10)
@@
$\sum_{j=0}^{n-1}f_{s}\left(\frac{j+\theta}{n}\right)$ $=n\int_{0}^{1}f_{s}(x)\mathrm{d}x+\sum_{k=0}^{l}a_{k}(s)(\zeta(2(s-k),\theta)+\zeta(2(s-k),1-\theta))\left(\frac{2\pi}{n}\right)^{2(k-s)}$
$+\mathcal{O}(n^{2(s-l-1)}).$
@@
$+\underset{k\neq s}{\sum_{0\leq k\leq l}}b_{k}(s)(\zeta(1+2(s-k),\theta)-\zeta(1+2(s-k),1-\theta))\left(\frac{2\pi}{n}\right)^{2(k-s)-1}$
$+\mathcal{O}(n^{1+2(s-l-1)}),$
@@
$\sum_{j=0}^{n-1}g_{s}\left(\frac{j+\theta}{n}\right)$ $=-\frac{n}{2\pi s}\Bigg[\sum_{k=0}^{l}b_{k}(s)\frac{s}{2(s-k)}\partial_{\theta}\zeta_{\mathbb{R}/\mathbb{Z}}(s-k,\theta)n^{2(s-k)}$
$+\mathcal{O}(n^{2(s-l-1)})\Bigg],$